To make 500 liters of a 25% acid solution. How many liters of 30% and18% acid solutions are needed to fill the order?
(Linear Mixture problem)

Answer by mananth(16946)

let the quantity of 30 % acid required be x liters
the 18% acide required will be 500 - x liters
we need 500 liters 25% acid solution
0.3x+0.18(500-x) = 0.25 * 500
0.3x+90-0.18x=125
0.12x=35
x= 35/0.12
291.66 liters of 30% acid is required.
balance will be 18% acid